# Multi-agent cooperation through learning-aware policy gradients

### v1

@@ Appendix C Proofs
$\nabla_{\phi^{i}}\bar{J}(\phi^{i})=\mathbb{E}_{\bar{P}^{\phi^{i}}}\left[\sum_{b=1}^{B}\sum_{l=1}^{MT}\nabla_{\phi^{i}}\log\pi^{i}(a^{i,b}_{l}\mid h^{i,b}_{l})\left(\frac{1}{B}\sum_{b^{\prime}=1}^{B}\sum_{k=l}^{MT}R_{k}^{i,b^{\prime}}\right)\right].$ (19)
@@
Applying the policy gradient theorem (Sutton et al., 1999) to the batched co-player shaping POMDP, for the minibatch-joint policy $\bar{\pi}^{i}(\bar{a}_{l}^{i}\mid\bar{h}_{l}^{i};\phi^{i})$ leads to
$\nabla_{\phi^{i}}\bar{J}(\phi^{i})=\mathbb{E}_{\bar{P}^{\phi^{i}}}\left[\sum_{l=1}^{MT}\nabla_{\phi^{i}}\log\bar{\pi}^{i}(\bar{a}_{l}^{i}\mid\bar{h}_{l}^{i};\phi^{i})\left(\frac{1}{B}\sum_{b^{\prime}=1}^{B}\sum_{k=l}^{MT}R_{k}^{i,b^{\prime}}\right)\right]$
Factorizing the minibatch-joint policy $\bar{\pi}$ as $\bar{\pi}^{i}(\bar{a}^{i}_{l}\mid\bar{h}^{i}_{l};\phi^{i})=\prod_{b=1}^{B}\pi^{i}(a^{i,b}_{l}\mid h^{i,b}_{l})$ leads to
$\nabla_{\phi^{i}}\bar{J}(\phi^{i})=\mathbb{E}_{\bar{P}^{\phi^{i}}}\left[\sum_{b=1}^{B}\sum_{l=1}^{MT}\nabla_{\phi^{i}}\log\pi^{i}(a^{i,b}_{l}\mid h^{i,b}_{l})\left(\frac{1}{B}\sum_{b^{\prime}=1}^{B}\sum_{k=l}^{MT}R_{k}^{i,b^{\prime}}\right)\right]$
thereby concluding the proof. ∎
Theorem C.2.
@@
Appendix D Deriving prior shaping algorithms from Eq. 1
@@
Appendix E Detail on baseline learning rules
E.1 Batch-unaware COALA PG
@@
^$\hat{\nabla}_{\phi^{i},\text{batch-unaware}}\bar{J}(\phi^{i})=\mathbb{E}_{\bar{P}^{\phi^{i}}}\left[\frac{1}{B}\sum_{b=1}^{B}\sum_{l=1}^{MT}\nabla_{\phi^{i}}\log\pi^{i}(a^{i,b}l\mid h^{i,b}_{l})\sum_{k=l}^{MT}R_{k}^{i,b}\right].$ (23)
E.2 M-FOS
@@
^$\hat{\nabla}_{\phi^{i},\text{M-FOS}}\bar{J}(\phi^{i})=\mathbb{E}_{\bar{P}^{\phi^{i}}}\left[\sum_{b=1}^{B}\sum_{l=1}^{MT}\nabla_{\phi^{i}}\log\pi^{i}(a^{i,b}_{l}\mid h^{i,b}_{l})\left(\sum_{k=l}^{Tm_{l}}R_{k}^{i,b}+\frac{1}{B}\sum_{b^{\prime}=1}^{B}\sum_{k=Tm_{l}+1}^{MT}R_{k}^{i,b^{\prime}}\right)\right].$ (24)
@@
Appendix F Software

### v2

@@ Appendix D Proofs
$\nabla_{\phi^{i}}\bar{J}(\phi^{i})=\mathbb{E}_{\bar{P}^{\phi^{i}}}\left[\sum_{b=1}^{B}\sum_{l=1}^{MT}\nabla_{\phi}\log\pi^{i}(a^{i,b}_{l}\mid h^{i,b}_{l})\left(\frac{1}{B}\sum_{l^{\prime}=l}^{m_{l}T}r_{l^{\prime}}^{i,b}+\frac{1}{B}\sum_{b^{\prime}=1}^{B}\sum_{l^{\prime}=m_{l}T+1}^{MT}r_{l^{\prime}}^{i,b^{\prime}}\right)\right].$ (19)
@@
In the co-player shaping batched POMDP there is only one agent that is relevant for the policy gradient, as all other agents are naive learners and subsumed in the environment dynamics. Hence, to avoid overloading notations, we drop the $i$ superscript in the parameters, actions, policy and histories. Furthermore, we use the notation $a_{l}=\{a_{l}^{b}\}_{b=1}^{B}$ and similarly for $h_{l}$.
We start by writing down the gradient of $\bar{J}(\phi)=\mathbb{E}_{\bar{P}^{\phi^{i}}}\left[\frac{1}{B}\sum_{b=1}^{B}\sum_{l=0}^{MT}R^{i,b}_{l}\right]$, making the summations in its expectation explicit.
$\nabla_{\phi}\bar{J}(\phi)=\nabla_{\phi}\sum_{l=0}^{MT}\sum_{r_{l}\in\bar{\mathcal{R}}}\sum_{a_{l}\in\bar{\mathcal{A}}}\sum_{h_{l}\in\bar{\mathcal{H}}_{l}}\bar{P}^{\phi}(h_{l})\bar{\pi}(a_{l}\mid h_{l})\bar{P}(r_{l}\mid h_{l},a_{l})\frac{1}{B}\sum_{b}r_{l}^{b}$
with ¯=×$\bar{\mathcal{R}}=\times_{b}\mathcal{R}$ the joint reward space, ¯$\bar{\mathcal{H}}_{l}$ the joint space over possible batched histories up until timestep $l$, and $\bar{\pi}(a_{l}\mid h_{l})=\prod_{b=1}^{B}\pi(a_{l}^{b}\mid h_{l}^{b})$. Applying the chain rule leads to
$\nabla_{\phi}\bar{J}(\phi)=$ $\sum_{l=0}^{MT}\sum_{r_{l}\in\bar{\mathcal{R}}}\sum_{a_{l}\in\bar{\mathcal{A}}}\sum_{h_{l}\in\bar{\mathcal{H}}_{l}}\nabla_{\phi}\bar{P}^{\phi}(h_{l})\bar{\pi}(a_{l}\mid h_{l})\bar{P}(r_{l}\mid h_{l},a_{l})\frac{1}{B}\sum_{b}r_{l}^{b}\ldots$
$+\sum_{l=0}^{MT}\sum_{r_{l}\in\bar{\mathcal{R}}}\sum_{a_{l}\in\bar{\mathcal{A}}}\sum_{h_{l}\in\bar{\mathcal{H}}_{l}}\bar{P}^{\phi}(h_{l})\nabla_{\phi}\bar{\pi}(a_{l}\mid h_{l})\bar{P}(r_{l}\mid h_{l},a_{l})\frac{1}{B}\sum_{b}r_{l}^{b}$
as the reward dynamics $\bar{P}(r_{l}\mid h_{l})$ are independent of the policy parameterization $\phi$. We first investigate the gradient of the marginal distribution $\nabla_{\phi}\bar{P}^{\phi}(h_{l})$, by marginalizing over the joint trajectory distribution.
$\nabla_{\phi}\bar{P}^{\phi}(h_{l})=\nabla_{\phi}\sum_{\{a_{l^{\prime}}\in\bar{\mathcal{A}},\{h_{l^{\prime}}\in\bar{\mathcal{H}}_{l^{\prime}}\}_{l^{\prime}<l}}\bar{P}(h_{l}\mid h_{l-1},a_{l-1})\prod_{l^{\prime}<l}\bar{P}(h_{l^{\prime}}\mid h_{l^{\prime}-1},a_{l^{\prime}-1})\bar{\pi}(a_{l^{\prime}}\mid h_{l^{\prime}})$
=$=\sum_{\{a_{l^{\prime}}\in\bar{\mathcal{A}},\{h_{l^{\prime}}\in\bar{\mathcal{H}}_{l^{\prime}}\}_{l^{\prime}<l}}\bar{P}(h_{l}\mid h_{l-1},a_{l-1})\prod_{l^{\prime}<l}\bar{P}(h_{l^{\prime}}\mid h_{l^{\prime}-1},a_{l^{\prime}-1})\bar{\pi}(a_{l^{\prime}}\mid h_{l^{\prime}})\sum_{l^{\prime\prime}<l}\nabla_{\phi}\log\bar{\pi}(a_{l^{\prime\prime}}\mid h_{l^{\prime\prime}})$
$=\mathbb{E}_{\bar{P}^{\phi}}\left[\sum_{l^{\prime}<l}\nabla_{\phi}\log\bar{\pi}(a_{l^{\prime}}\mid h_{l^{\prime}})\right]$
with $\{a_{l^{\prime}}\in\bar{\mathcal{A}},\{h_{l^{\prime}}\in\bar{\mathcal{H}}_{l^{\prime}}\}_{l^{\prime}<l}$ the joint space over all actions and histories over timesteps $l^{\prime}<l$. In the second line, we used the chain rule and $\nabla_{\phi}\bar{\pi}=\bar{\pi}\nabla_{\phi}\log\bar{\pi}$. In the third line we renamed the index $l^{\prime\prime}$ to $l^{\prime}$.
Filling this expression for $\nabla_{\phi}\bar{P}^{\phi}(h_{l})$ into our expression for $\nabla_{\phi}\bar{J}(\phi)$, combined with the log trick $\nabla_{\phi}\bar{\pi}=\bar{\pi}\nabla_{\phi}\log\bar{\pi}$ and using the expectation notation for clarity of notation, we end up with
$\nabla_{\phi}\bar{J}(\phi)=\mathbb{E}_{\bar{P}^{\phi}}\left[\frac{1}{B}\sum_{b=1}^{B}\sum_{l=0}^{MT}r_{l}^{b}\sum_{l^{\prime}\leq l}\nabla_{\phi}\log\bar{\pi}(a_{l^{\prime}}\mid h_{l^{\prime}})\right]$
Reordering the summations, and using $\bar{\pi}(a_{l}\mid h_{l})=\prod_{b=1}^{B}\pi(a_{l}^{b}\mid h_{l}^{b})$ leads to
$\nabla_{\phi}\bar{J}(\phi)=\mathbb{E}_{\bar{P}^{\phi}}\left[\sum_{b=1}^{B}\sum_{l=1}^{MT}\nabla_{\phi}\log\pi(a^{b}_{l}\mid h^{b}_{l})\left(\frac{1}{B}\sum_{b^{\prime}=1}^{B}\sum_{l^{\prime}=l}^{MT}r_{l^{\prime}}^{b^{\prime}}\right)\right].$
Finally, actions can only influence the other parallel trajectories through the parameter updates of the naive learners in the environment, which takes place at the inner episode boundaries. Hence, during the current inner episode (before a naive learner update takes place), rewards $r_{l}^{b^{\prime}}$ are independent of actions $a_{l}^{b}$ for $b\neq b^{\prime}$. As $\mathbb{E}_{a\sim\pi}\left[c\log\pi(a\mid h)\right]=0$ for a constant $c$ independent of the actions, the policy gradient is equal to
$\nabla_{\phi}\bar{J}(\phi)=\mathbb{E}_{\bar{P}^{\phi}}\left[\sum_{b=1}^{B}\sum_{l=1}^{MT}\nabla_{\phi}\log\pi(a^{b}_{l}\mid h^{b}_{l})\left(\frac{1}{B}\sum_{l^{\prime}=l}^{m_{l}T}r_{l^{\prime}}^{b}+\frac{1}{B}\sum_{b^{\prime}=1}^{B}\sum_{l^{\prime}=m_{l}T+1}^{MT}r_{l^{\prime}}^{b^{\prime}}\right)\right].$
with $m_{l}$ the inner episode index corresponding to timestep $l$, thereby concluding the proof.
Theorem D.2.
@@
Appendix E Relating COALA-PG to Learning with Opponent-Learning Awareness (LOLA)
In this section, we establish a formal relationship between COALA-PG and Learning with Opponent-Learning Awareness (LOLA; Foerster et al., 2018a), the seminal work that spearheaded the learning awareness field. In doing so, we further show how COALA-PG can be used to derive a new LOLA gradient estimator that does not require higher-order derivative estimates.
LOLA considers a POSG (c.f. Section 2.2) with agent policies $(\phi^{i},\phi^{-i})$, and expected return $J^{i}(\phi^{i},\phi^{-i})$. Recall from Section 4 that instead of estimating the naive gradients $\nabla_{\phi^{i}}J^{i}(\phi^{i},\phi^{-i})$, LOLA anticipates that co-players update their parameters with $M$ naive gradient steps. The improved LOLA-DICE (Foerster et al., 2018b) update reads:
with $\frac{\mathrm{d}}{\mathrm{d}\phi^{i}}$ the total derivative taking into account the effect of $\phi^{i}$ on the parameter updates $\Delta_{q}\phi^{-i}$, and $\frac{\partial}{\partial\phi^{-i}}$ the partial derivative.
Despite the apparent dissimilarities of the two algorithms, we now show in Theorem E.1 that as a special case of COALA-PG, we can estimate the gradient of a similar objective to the LOLA-DICE method. We consider a mixed agent group of meta agents $(\phi^{i},\phi^{-i})$, and naive learners.
Theorem E.1.
Appendix F Deriving prior shaping algorithms from Eq. 1
@@ Appendix E Relating COALA-PG to Learning with Opponent-Learning Awareness (LOLA)
Appendix G Detail on baseline methods
G.1 Batch-unaware COALA PG
@@
^$\hat{\nabla}_{\phi^{i}}^{\text{batch-unaware}}\bar{J}(\phi^{i})=\mathbb{E}_{\bar{P}^{\phi^{i}}}\left[\frac{1}{B}\sum_{b=1}^{B}\sum_{l=1}^{MT}\nabla_{\phi^{i}}\log\pi^{i}(a^{i,b}l\mid h^{i,b}_{l})\sum_{k=l}^{MT}R_{k}^{i,b}\right].$ (25)
Figure 11 visualizes the main difference between the COALA-PG update and the batch-unaware COALA-PG update.
Figure 11: Diagram visualizing the difference between the COALA-PG update and the batch unaware COALA policy gradient update.
G.2 M-FOS
@@
^$\hat{\nabla}_{\phi^{i}}^{\text{M-FOS}}\bar{J}(\phi^{i})=\mathbb{E}_{\bar{P}^{\phi^{i}}}\left[\sum_{b=1}^{B}\sum_{l=1}^{MT}\nabla_{\phi^{i}}\log\pi^{i}(a^{i,b}_{l}\mid h^{i,b}_{l})\left(\sum_{k=l}^{Tm_{l}}R_{k}^{i,b}+\frac{1}{B}\sum_{b^{\prime}=1}^{B}\sum_{k=Tm_{l}+1}^{MT}R_{k}^{i,b^{\prime}}\right)\right].$ (26)
@@
Appendix H Additional experimental results
H.1 Balancing gradient contributions
The main difference between COALA-PG (c.f. Eq. 3) and M-FOS is the scaling of the current reward-episode. The main difference between COALA-PG equation 3 and M-FOS equation 7 is that COALA-PG scales the return of the current inner episode by $\frac{1}{B}$, whereas M-FOS does not. This scaling is crucial, as the the influence of an action on future inner episodes is of order $O(\frac{1}{B})$ because the naive learner averages its update over the $B$ inner episode trajectories. Hence, by scaling the inner episode return by $\frac{1}{B}$, COALA-PG ensures that both contributions have the same scaling w.r.t. $B$, and finally by summing the policy gradient over the minibatch instead of averaging, we end up with a policy gradient of $O(1)$. M-FOS and batch-unaware COALA do not scale the current inner episode return, which causes the co-player shaping learning signal to vanish w.r.t. the current inner episode return, resulting in poor shaping performance.
Figure 12: Y-axis: Ratio of the magnitude of the policy gradient contribution arising from future inner episode returns (the co-player shaping learning signal) w.r.t. the gradient contribution arising from the current inner episode return. X-axis: meta gradient steps on the iterated prisoner’s dilemma, in a mixed group setting corresponding to the setting of Fig. 5C, and an increased meta-batch size of 2048 to reduce the variance of the gradient estimates.
Figure 12 confirms that empirically, COALA-PG correctly balances the gradient contributions from the current inner episode return with those from future inner episode returns. In contrast, M-FOS and batch-unaware COALA have unbalanced gradient contributions, with the co-player shaping gradient contribution vanishing w.r.t. the current inner episode return contribution.
H.2 Detailed results on CleanUp-lite
Figure 13: Agents trained by COALA-PG against naive agents only successfully shape them in CleanUp-lite. (A) COALA-PG-trained agents better shape naive opponents compared to baselines, obtaining higher return. (B and C) Analyzing behavior within a single meta-episode after training reveals that COALA outperforms baselines and shapes naive agents, (i) exhibiting a lower cleaning discrepancy (absolute difference in average cleaning time between the two agents), and (ii) being less often zapped. (D) Average reward of naive learners is higher when playing with COALA-PG agents compared to other agents. (E and F): COALA-PG results in lower average pollution level and higher average apple level. Shaded regions indicate standard deviation computed over 5 seeds.
Figure 14: Agents trained with COALA-PG against a mixture of naive and other meta agents learn to cooperate in CleanUp-lite. (A) COALA-PG-trained agents obtain higher average reward than baseline agents when playing against each other. (B and C): COALA-PG leads to a more fair division of cleaning efforts and lower zapping rates. (D and E): COALA-PG results in lower average pollution level and higher average apple level. Shaded regions indicate standard deviation.
H.3 Training a COALA-PG agent versus an M-FOS agent on iterated prisoner’s dilemma
We investigate training a COALA-PG versus an M-FOS agent, our strongest performing meta-agent baseline. In this setup, we still use a mixture of naive learners and meta agents. Only now when a COALA-PG agent either plays against an M-FOS agent or naive learner, and an M-FOS agent either plays against a COALA-PG agent or a naive learner. We found that COALA-PG successfully shapes M-FOS into cooperation, reaching average rewards of $0.850\pm 0.037$ for COALA-PG and $0.853\pm 0.017$ for M-FOS. This is in stark contrast to when M-FOS agents only play against other M-FOS agents and naive learners, which converges to mutual defection.
We did not investigate training a COALA-PG agent versus a LOLA agent, as the training setup of both agents is fundamentally different, and would pose an unfair disadvantage for LOLA agents. A LOLA agent learns on the same timescale as naive learners, taking a few look-ahead steps into account in its update. Hence, the full learning trajectory of a LOLA agent is considered as one meta trajectory for a COALA-PG agent. As COALA-PG agents would play many different meta trajectories, always against a freshly initialized LOLA agent, this would give a COALA-PG detailed information about the learning behavior of LOLA agents, whereas LOLA agents are left in the dark as they cannot observe the meta updates of COALA-PG. Hence, COALA-PG would extort LOLA agents similar to naive learners.
Appendix I Software